INAF queue observing, 2014 Jan 29-30

Observer: E. Sani, F. Cusano, S. Bisogni, R. Carini
Telescope Operator: D. Gonzales Huerta
Support Astronomer: O. Kuhn
Instrument: MODS

Summary

During the twilight we have thin cirrus passing, wind at 15 m/s and 40% humidity .
N.B. The seeing from the GCS is overestimated by 10% because the MODs guider is out of focus.
At the beginning of science integration the seeing is 1.4 on the GCS and 1.6 on the DIMM amd thin clouds are passing.
Thick clouds passing during the transit. The WFS was not sending corrections to the primary and observations are affected by astigmatism. The SA calls J. C. Guerra to try to increase the WFS exp time, but this is coupled with the guider. To do not miss the guiding we give it up.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard G191b2b dual grating - airmass 1.1
Nascimbeni_DDT GJ3470b dual grating 4h 40 m Executed with some issues
Gilli Q1 J1048 Q1 red grating 1h 30m Executed out of seeing specs
Piconcelli C1 RBS 1273 red grating 5 m Executed out of seeing specs
Piconcelli C2 RBS 1124 red grating 5 m Executed out of seeing specs

Details

00:44 The TO opens the dome

01:45 Slew to the standard G191b2b
  • acqMODS g191b2b_ls.acq mods1r.20140129.0010-0014
  • modsAlign -r mods1r 0011

02:01 execMods g191b2b_ls.obs mods1r.20140129.0015 mods1b 0010
  • the average seeing on the GCS is 1.2

02:17 We slew to Nascimbeni_DDT GJ3470b transit. Too early

02:20 Slew back to the standard to integrate also the redgrating only. Good for Gilli_04 and Piconcelli programs.
  • acqMODS g191b2b_ls.acq mods1r.20140129.0016-0019
  • modsAlign -r mods1rr.20140129.0017

02:34 execMods g191b2b_ls.obs mods1r0r.20140129.020

02:43 We slew to Nascimbeni_DDT GJ3470b transit.
  • acqMODS mods.2.GJ3470.acq mods1rr.20140129.0021-0024
  • modsAlign mods.2.GJ3470.mms mods1rr.20140129.0021 mods1r 0022

03:04 execMods mods.2.GJ3470.obs mods1r0r.20140129.011-0103 mods1b 0025-0058
  • at the beginning of integration the seeing is 1.4 on the GCS and 1.6 on the DIMM. Thin clouds are passing.
  • Thin clouds passing during the integration.
  • Due to thick clouds passing the WFS is no able to send corrections to collimate the primary. This is producing astigmatism. We thus commit a 0.0001 offset in the y direction to force the probe.
  • The SA calls J. C. Guerra to know if it is possible to increase the WFS exposure time independently from the guider exposure time. This is not possible.
  • Thick clouds lasted for ~1.5h after this the WFS starts to send corrections.
  • Stuck of the red side reading and writing data on /newdata. We do have to start exposure manually. We lost image mods1r 0085 because the detector flush lasted too long and was not possible to write the fits.

07:41 We stop the script as required

07:43 We slew to Gilli_Q1
  • acqMODS 4Q1_J1048_M1_mos.acq mods1r.20140129.0104-0107
  • seeing 1.6

08:08 execMODS 4Q1_J1048_M1_mos_4.obs mods1r.20140130.0001-0002
  • update the name of the files to UT date 20140130, we forgot at the beginning of the night to do the getpudate from the MODS dashboard
  • seeing 1.8

08:44 execMODS 4Q1_J1048_M1_mos_5.obs mods1r.20140130.0003
  • very bad seeing during the integration. Data out of specs.

09:12 Strong wind. The TO closes the dome

09:41 We start calibrations
  • execMODS mos.2.GJ340 mods1r 0004-0008 mods1b 0001-0005 SATURATED Flats + Lamps with some lines saturated

10:05 Dome opened

10:07 acqMODS 4Q1_J1048_M1_mos.acq mods1r 0009-0011
  • modsAlign mods.6.mms mods1r 0009-0010

10:40 execMODS 4Q1_J1048_M1_mos_4.obs mods1r 0012-0013
  • seeing 1.2-2

11:15 execMODS 4Q1_J1048_M1_mos_4.obs mods1r 0014
  • seeing 2.

11:40 Error No offset sent: the probe lost the guiding star. This was due to exceptionally bad seeing.
  • the TO slew to a bright star to collimate. The guider can not

12:15 Slew to piconcelli_C1 RBS 1273 mods1r 0015-0018
  • acqMODS C1_LS.acq
  • modsAlign -l mods1r 0015-0016

12:40 execMODS C1_LS.obs mods1r 0019-0020
  • seeing 1.4-1.5

12:54 Slew to piconcelli_C2 RBS 1124 mods1r 0021-0024
  • acqMODS C1_LS.acq
  • modsAlign -l -B mods1r 0021-0022

13:12 execMODS C2_LS.acq mods1r 0025-0026
  • seeing 1.4

13:24 The TO closes the Dome
  • execMODS mods.2.GJ3470_flats.cal mods1r 0027-0029 mods1b 0006-0007 all saturated but mods1b.0006
  • execMODS mods.2.GJ3470_flats.cal mods1r 0030-0032 mods1b 0008-0009 GOOD Flats
  • execMODS mods.2.GJ3470_lamp.cal mods1r 0033-0034 mods1b 0010-0011 GOOD Lamps
  • execMODS mods.J1048_M2_mos.cal mods1r 0036-0039 GOOD Flat+Lamps
  • execMODS 39A_flat.cal mods1r 0040-0048 mods1b 0012-0022 GOOD Flats

-- EleanoraSani - 30 Jan 2014
Topic revision: r7 - 05 Feb 2014, EleanoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback